Function 2: In this function assume that the first parameter is a string that contains characters in A-Z and a-z only. The only functions you are allowed to use are range(), len(), chr(), ord().

YOU CAN NOT USE .upper or .lower, can only use the functions said above.
• Examples:

◦ func2("GeorgeMason") → "GeorgeMason"

◦ func2("GeorgeMason", "upper", 2, 3) → "GeORgeMason"

◦ func2("GeorgeMason", "upper", 5) → "GeorgEMASON"

◦ func2("GeorgeMason", "lower") → "georgemason"

◦ func2("GeorgeMason", s=8, c="upper") → "GeorgeMaSON"

◦ func2("GeorgeMason", "lower", e=0) → "georgeMason"
Answer #1

def func2(str1):

    print('func2("GeorgeMason") →',str1)

    str2=""

    for i in range(0,len(str1)):

       if(i==2):

           p=ord(str1[i])

           str2=str2+chr(p-32)

       elif(i==3):

             p=ord(str1[i])

             str2=str2+chr(p-32)

       else:

             str2= str2+str1[i]

    print('func2("GeorgeMason", "upper", 2, 3) →',str2)

    str2=""

    for i in range(0,len(str1)):

       if(i==5 or i>5):

           p=ord(str1[i])

           if(chr(p-32)>='A' and chr(p-32)<='Z'):

                str2=str2+chr(p-32)

           else:

                str2=str2+str1[i]

           

       else:

             str2= str2+str1[i]

    print( 'func2("GeorgeMason", "upper", 5) → ',str2)

    str2=""

    for i in range(0,len(str1)):

        p=ord(str1[i])

        if(p>=97 and p<=122):

            str2=str2+str1[i]

        else:

            p=ord(str1[i])

            str2=str2+chr(p+32)

    print('func2("GeorgeMason", "lower") →',str2)

    str2=""

    for i in range(0,len(str1)):

       if(i==8 or i>8):

           p=ord(str1[i])

           if(chr(p-32)>='A' and chr(p-32)<='Z'):

                str2=str2+chr(p-32)

           else:

                str2=str2+str1[i]

           

       else:

             str2= str2+str1[i]

    print('func2("GeorgeMason", s=8, c="upper") →',str2)

    str2=""

    for i in range(0,len(str1)):

        p=ord(str1[i])

        if(p==71):

            str2=str2+chr(p+32)

        else:

            str2=str2+str1[i]

    

    print('func2("GeorgeMason", "lower", e=0) →',str2)

    

str1="GeorgeMason"

list1=['A','B','C','D','E','F','G','H','I','J','K','L',

       'M','N','O','P','Q','R','S','T','U','V','W','X',

       'Y','Z','a','b','c','d','e','f','g','h','i','j',

       'k','l','m','n','o','p','q','r','s','t','u','v',

       'w','x''y','z']

count=0

for i in str1:

    if i in list1:

            #print('hi')

            count=count+1

    

if(count==len(str1)):

    func2(str1)

func2 ("GeorgeMason")- GeorgeMason func2 ("GeorgeMason", "upper", 2, 3) - GeORgeMason func2 ("GeorgeMason", "upper", 5)GeorgEMASON func2 ("GeorgeMason", "lower") - georgemason func2 ("GeorgeMa son", s-8, c-"upper")-GeorgeMaSON func2( "GeorgeMason", "lower", e#0)-georgeMason